# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v4

@@ 6 Discussion
In this paper, we put forward an economical framework to design Bayesian A/B tests involving multiple metrics while satisfying criteria for the FDR and average power. This framework recommends optimal sample sizes and decision thresholds that are tailored to the specified dependence structure between the hypothesis tests and accommodates uncertainty regarding which hypotheses are false. The efficiency of this framework stems from considering a proxy for the joint sampling distribution of posterior probabilities based on large-sample theory to motivate estimating the true sampling distributions at only two sample sizes. This approach significantly decreases the computational overhead required to design Bayesian A/B tests, making them much more attractive and accessible to practitioners that want to control the FDR.
The methods proposed in this paper could be extended in various aspects. Further work could develop more sophisticated and efficient optimization schemes to determine decision thresholds $\{\gamma_{k}\}_{k=1}^{K}$ that control the FDR. It may also be useful to consider design criteria for A/B tests other than the FDR or the FWER. This consideration could be framed using the three types of metrics defined by Kohavi et al. (2020) based on the evidence supporting metric improvement or deterioration in group B. In particular, positive and negative metrics respectively have statistically significant evidence of improvement and deterioration, and flat metrics have no statistically significant evidence of improvement or deterioration. Kohavi et al. (2020) recommended implementing a new online experience if all metrics are flat or positive, with at least one metric being positive. To classify the $k^{\text{th}}$ metric as positive, negative, or flat, we would need to compare $Pr(H_{1,k}~{}|~{}data)$ to both lower and upper decision thresholds in an alternative design framework.
@@
These materials include a detailed description of the conditions for Theorem 1 and its proof, as well as additional context for the example in Section 5. The code to conduct the numerical studies in the paper is available online: https://github.com/lmhagar/ABTestFDR.

### v5

@@ 6 Discussion
In this paper, we put forward an economical framework to design Bayesian A/B tests involving multiple metrics while satisfying criteria for the Bayesian FDR and average power. This framework recommends optimal sample sizes and decision thresholds that are tailored to the specified dependence structure between the hypothesis tests and accommodates uncertainty regarding which hypotheses are false. The efficiency of this framework stems from considering a proxy for the joint sampling distribution of posterior probabilities based on large-sample theory to motivate estimating the true sampling distributions at only two sample sizes. This approach significantly decreases the computational overhead required to design Bayesian A/B tests, making them much more attractive and accessible to practitioners that want to control the Bayesian FDR.
The methods proposed in this paper could be extended in various aspects. Further work could develop more sophisticated and efficient optimization schemes to determine decision thresholds $\{\gamma_{k}\}_{k=1}^{K}$ that control the Bayesian FDR. It may also be useful to consider design criteria for A/B tests other than the Bayesian FDR or the FWER. This consideration could be framed using the three types of metrics defined by Kohavi et al. (2020) based on the evidence supporting metric improvement or deterioration in group B. In particular, positive and negative metrics respectively have statistically significant evidence of improvement and deterioration, and flat metrics have no statistically significant evidence of improvement or deterioration. Kohavi et al. (2020) recommended implementing a new online experience if all metrics are flat or positive, with at least one metric being positive. To classify the $k^{\text{th}}$ metric as positive, negative, or flat, we would need to compare $Pr(H_{1,k}~|~data)$ to both lower and upper decision thresholds in an alternative design framework.
@@
These materials include a detailed description of the conditions for Theorem 1, the proofs of Theorem 1 and Corollary 1, and additional context for the example in Section 5. The code to implement our methods and conduct our numerical studies is available online: https://github.com/lmhagar/ABTestFDR.